# Adaptive Multilevel Splitting: First Application to Rare Event Derivative Pricing

### v2

@@ 1 Introduction
The accurate and efficient pricing of financial derivatives is increasingly critical in modern markets, where advanced numerical methods are required for complex instruments [jrfm12010035]. The computational challenges of rare event simulation extend beyond academic interest, creating bottlenecks that affect market functionality. Inaccurate pricing of low probability events limits the ability of market makers to provide competitive quotes, reducing liquidity for these instruments [Muellerleile_2025]. This difficulty is pronounced in the insurance sector, where parametric products depend on binary triggers linked to observable parameters such as earthquake magnitude or wind speed [undici, LARSSON2023100345]. Computational limitations restrict coverage of catastrophic risks and constrain the development of innovative risk transfer mechanisms in financial and insurance markets.
@@
These challenges are evident in binary options, which share structural similarities with parametric insurance through trigger-based payoffs. Their discontinuous structure pays a fixed amount if the underlying asset crosses a predetermined barrier at expiration and zero otherwise [shreve2004stochastic2, shreve2004stochastic1]. This all-or-nothing feature makes pricing highly sensitive to the probability of rare events, particularly for deep out-of-the-money contracts where accurate tail estimation is critical.
@@
Addressing these difficulties naturally leads to simulation-based techniques. Monte Carlo methods are widely used for pricing complex derivatives due to their flexibility in high-dimensional settings [glasserman2004monte]. The convergence rate of $O(N^{-1/2})$ creates a computational bottleneck, especially for binary options with low exercise probabilities. Reliable estimation in such cases typically requires millions of paths, rendering crude Monte Carlo impractical [beck2015rareeventsimulation, bucklew2004introduction].
@@
Classical variance reduction techniques attempt to address these challenges. Antithetic variates reduce variance through negative correlation between paired samples [hammersley1956new, glasserman2004monte], but the theoretical gain is bounded by a factor of two [varredu]. Control variates can be more effective but require auxiliary variables that are both analytically tractable and highly correlated with the target payoff [rasmussen2005control]. For discontinuous payoffs such as binary options, such variables are difficult to construct, limiting applicability.
@@
More advanced methods have been developed. Importance sampling modifies the probability measure to increase the frequency of rare outcomes and applies likelihood ratio weighting to remove bias [glasserman2004monte, imps]. Its effectiveness depends on the design of suitable distributions, which is problem-specific and difficult to generalize [swiler2010importance]. Another prominent approach is multilevel Monte Carlo (MLMC), which reduces complexity by combining simulations on coarse and fine discretizations [m1, m2]. While efficient for path-dependent derivatives, MLMC is not tailored to extreme event pricing, focusing instead on reducing overall cost.
@@
Recent research combines these techniques to overcome individual limitations. Hybrid methods integrate MLMC with importance sampling to improve efficiency while concentrating sampling in critical regions [BenAlaya17022023, kebaier2018coupling]. Machine learning further enhances importance sampling, with neural networks learning tilting parameters [muller2019neural] and tensor train decompositions enabling high-dimensional distribution approximation [cui2024deep].
@@
This work addresses the computational challenges of binary option pricing by applying the adaptive multilevel splitting (AMS) method [Cérou27022007]. AMS extends classical splitting techniques for rare event simulation [garvels2000splitting] and builds on the foundations of sequential Monte Carlo [doucet2001sequential]. Originally developed in reliability analysis and statistical physics [c1, baars2021application, innes2024adaptive, refId0], AMS decomposes a rare event into a sequence of more frequent conditional events, transforming a single intractable estimation into multiple tractable subproblems. Although AMS has achieved strong results in other scientific domains, no prior applications are documented in financial derivatives pricing. Recent advances provide theoretical guarantees of unbiasedness and convergence [brehier2016unbiasedness, cerou2016fluctuation, cerou2023adaptive], creating a rigorous basis for its use in finance.
@@
Within quantitative finance, interacting particle systems and Sequential Monte Carlo (SMC) methods have already been applied to derivative pricing, most notably in the barrier-option framework of [Shevchenko_2016]. In that setting, the option price is expressed through a Feynman–Kac representation whose potential functions enforce survival within the barrier region, and resampling is used to maintain a population of feasible trajectories. Numerical stability for rare events is typically achieved by tempering or by carefully designing these potentials, while the intermediate “levels” that guide the particle system are fixed ex ante by the time discretisation or by a prescribed sequence of thresholds. Adaptive Multilevel Splitting (AMS) can be understood as a direct evolution of this idea [c1]: instead of resampling based on fixed survival sets or generic particle weights, AMS works with a scalar importance function on the whole trajectory and introduces an adaptive sequence of levels determined by the empirical distribution of the particles. At each iteration, a fixed fraction of the least-performing particles is killed and replaced by clones of better-performing ones, so that the intermediate levels separating typical from rare trajectories are constructed on the fly.
@@
The contributions of this study are fourfold. First, an AMS adaptation is introduced for binary option pricing under Black–Scholes and Heston dynamics [due, e0f45016-f730-320a-bedc-d34f406805b2], addressing the specific challenges of risk-neutral valuation and financial time series. Second, the sensitivity of the estimator to parameter choices, including the number of trajectories and resampling rates, is analyzed. Third, numerical experiments compare AMS to standard Monte Carlo, showing substantial gains for deep out-of-the-money options. Fourth, an open source Rcpp implementation is released, supporting Euler, Milstein, and Andersen discretizations [BallyTalay1996, HighamMaoSzpruch2013, Andersen2007], two importance functions, and six binary option variants, offering a flexible toolkit for rare event simulation in derivatives pricing. A fully documented and publicly available implementation is provided in the R package amsSim [mio], which has been released on CRAN and includes all the algorithms and numerical experiments presented in this paper.
@@ 2.1 Stochastic differential equation models
Numerical experiments are conducted under two standard models for asset price dynamics: the Black–Scholes model [due] and the Heston model [e0f45016-f730-320a-bedc-d34f406805b2]. These frameworks allow assessment of the robustness of the AMS approach across different model complexities.
For the Black–Scholes case the exact solution, obtained via logarithmic transformation, removes discretization error [BOYLE1977323, glasserman2004monte]:
@@
For the Heston model the variance process requires a scheme that preserves positivity and avoids bias. The quadratic–exponential (QE) method of Andersen [Andersen2007] is employed, the standard approach for accurate Heston simulation. It matches the first two conditional moments of $V_{t+\Delta t}\,|\,V_{t}$ and selects the update regime according to
@@
with $\epsilon\sim\mathcal{N}(0,1)$. The coefficients $\{K_{0},\dots,K_{4}\}$ and the parameters $a$, $b$, and $\psi$ are given explicitly in [Andersen2007].
@@ 2.3 Adaptive multilevel splitting (AMS)
Adaptive multilevel splitting (AMS) [Cérou27022007, c1] is a variance reduction method for estimating the probability of rare events. Instead of relying on brute–force Monte Carlo, which wastes almost all trajectories on paths that never approach the rare set, AMS iteratively concentrates simulation effort on trajectories that make progress towards the event of interest.
@@
which is sufficient to preserve unbiasedness [brehier2016unbiasedness]. Equality between the two sets is not required, although a score function that is more closely aligned with $D$ leads to a lower variance. For a complete trajectory $X=(X_{t})_{t\in[0,\tau_{f}]}$ we then define its score as
@@ 2.3.1 Theoretical properties of AMS
Let $X=(X_{t})_{t\geq 0}$ be a Markov process with importance function $\xi$ and rare set $D$. For fixed population size $N$ and killing parameter $K\in\{1,\dots,N-1\}$, AMS is well posed: at each iteration the cutting level $Z$ is an order statistic $Z=S^{(K)}$, and under standard regularity assumptions on $X$ (Feller property), on $\xi$ (continuity), and on the entrance into $D$ (strict entrance condition), the algorithm terminates almost surely after a finite number of iterations [c1].
@@
without requiring the stronger equivalence $\xi(x)\geq L_{\max}\iff x\in D$ [brehier2016unbiasedness, 10.1093/biomet/asy028, c1]. Unbiasedness extends to unnormalised measures
@@
where the asymptotic variance $\sigma^{2}(\varphi)$ can be characterised via the associated Fleming–Viot particle system [c1, doi:10.1137/18M1187477]. In particular, for $K=1$ and a rare event with probability $p$,
@@
While the unbiasedness of $\hat{p}_{\mathrm{AMS}}$ does not depend on the particular choice of the importance function $\xi$, the variance of the estimator is strongly affected by it. In particular, if $\xi$ is poorly adapted to the geometry of the rare event (for example, in analogy with a badly chosen importance sampling density), the resulting estimator may exhibit large variance and therefore a much slower convergence to the true value. In practice, variance is controlled by experimentally comparing alternative definitions of $\xi$ and by suitably adjusting the population size $N$ and the killing parameter $K$ [c1, brehier2016unbiasedness].
@@
AMS adapts intermediate levels and branching rates on the fly, removing the need for a priori specification as in classical Multilevel Splitting [Kahn1951SplittingParticleTransmission] or Sequential Monte Carlo [fba646e2e09d4f7f93e66f71554f16b7, cerou:inria-00071391]. The algorithm maintains a fixed population size $N$, ensuring robustness, parallel efficiency, and predictable memory use. It provides unbiased estimators for both rare event probabilities and unnormalised measures $\gamma(\varphi)$, enabling straightforward parallelization across independent runs [c1, brehier2016unbiasedness, 10.1093/biomet/asy028].
@@ 3.1 Antithetic variates: overview and limitations
Antithetic variates [gentle2009antithetic] reduce variance by pairing negatively correlated samples. In option pricing this corresponds to simulating each path together with its reflection obtained by negating Brownian increments. For monotone payoffs the estimator variance decreases, with a theoretical maximum reduction by a factor of two.
@@ 3.2 Control variates: overview and limitations
In practice, however, there is no unique, model–independent choice of $W$ that works well across products: an effective control variate must be engineered to closely mimic the structure of the specific payoff under the given dynamics. For example, for arithmetic Asian options one can use the optimized lower-bound proxy of [FusaiKyriakou2016]. This makes the technique very powerful when specialized, but at the cost of poor generalizability.
@@ 3.3 Multilevel Monte Carlo: overview and limitations
Multilevel Monte Carlo (MLMC) [m1, m2] exploits a hierarchy of approximations $X_{0},\dots,X_{L}$ of the same quantity. The telescoping identity
@@
reduces variance by coupling successive levels with shared randomness. The resulting estimator achieves mean square error $\mathcal{O}(\varepsilon^{2})$ at cost $\mathcal{O}(\varepsilon^{-2})$, compared to $\mathcal{O}(\varepsilon^{-3})$ for standard Monte Carlo [m1].
@@ 3.4 Importance sampling: overview and limitations
Importance sampling (IS) [tokdar2010importance] estimates $\psi=\mathbb{E}[h(X)]$ by sampling from an alternative density $g$ and reweighting:
@@
In rare event regimes IS becomes unstable. When exercise probabilities are of order $10^{-6}$, the equation $\psi^{\prime}(\theta)=a$ may lack a solution or yield extreme $\theta^{*}$, and evaluation of $e^{\theta Y}$ produces flat likelihood landscapes with sporadic spikes. In such cases Newton–Raphson and related solvers fail to converge, and stochastic optimisers are equally unreliable [casella2021choice].
@@
Variance explosion: an inappropriate choice of $g(y)$ can inflate the estimator’s variance instead of reducing it [tokdar2010importance].
Payoff specific design: effective importance sampling must be tailored to the payoff. Binary calls, binary puts, and Asian options require distinct tilting schemes, and multi-asset payoffs add combinatorial complexity [imps].
@@ 4.1 Importance function design
of [brehier2016unbiasedness, Theorem 3.2] holds. Then the AMS estimator of the risk–neutral probability $p=\mathbb{Q}(D)$ is
@@ 4.1.1 Heuristics and diagnostics for importance function design
From a theoretical perspective, a first requirement is that $\xi$ preserves the support of the event, in the sense that $x\in D\;\;\Longrightarrow\;\;\xi(x)\geq L_{\max},$ so that the sufficient condition of [brehier2016unbiasedness, Theorem 3.2] is satisfied. If this condition were violated, AMS could in principle miss part of $D$ and introduce bias. Even when unbiasedness holds, a poorly informative choice of $\xi$ (for instance, a score that is nearly uncorrelated with $\mathbf{1}_{D}$ or that does not increase as trajectories move towards $D$) inflates the variance and can lead to worse efficiency than crude Monte Carlo.
@@ 4.2 Impact of the selection parameter $K$ on algorithm performance
The selection parameter can be expressed either as the absolute number of discarded trajectories $K$, or equivalently as the discard fraction $k=K/N$, where $N$ is the population size. At each iteration exactly $K=kN$ particles (i.e., a fraction $k$ of the population) are killed and replaced by clones. Theoretical results show that, for an optimal importance function, the asymptotic variance is minimised in the limit $K=1$ [c1, doi:10.1137/18M1187477], but this regime is computationally impractical in realistic settings.
@@ 4.3 Impact of the number of trajectories N on algorithm performance
The particle count $N$ directly affects AMS performance. Larger $N$ reduces estimator variance but increases runtime due to higher simulation and sorting costs. Theoretical analysis shows complexity of order $N(-\log(p))\log(N)$, accounting for the sorting step and the generation of one new trajectory per iteration [c1, doi:10.1137/18M1187477].

### v3

@@ 1 Introduction
The accurate and efficient pricing of financial derivatives is increasingly critical in modern markets, where advanced numerical methods are required for complex instruments [53]. The computational challenges of rare event simulation extend beyond academic interest, creating bottlenecks that affect market functionality. Inaccurate pricing of low probability events limits the ability of market makers to provide competitive quotes, reducing liquidity for these instruments [41]. This difficulty is pronounced in the insurance sector, where parametric products depend on binary triggers linked to observable parameters such as earthquake magnitude or wind speed [43, 38]. Computational limitations restrict coverage of catastrophic risks and constrain the development of innovative risk transfer mechanisms in financial and insurance markets.
@@
These challenges are evident in binary options, which share structural similarities with parametric insurance through trigger-based payoffs. Their discontinuous structure pays a fixed amount if the underlying asset crosses a predetermined barrier at expiration and zero otherwise [49, 48]. This all-or-nothing feature makes pricing highly sensitive to the probability of rare events, particularly for deep out-of-the-money contracts where accurate tail estimation is critical.
@@
Addressing these difficulties naturally leads to simulation-based techniques. Monte Carlo methods are widely used for pricing complex derivatives due to their flexibility in high-dimensional settings [27]. The convergence rate of $O(N^{-1/2})$ creates a computational bottleneck, especially for binary options with low exercise probabilities. Reliable estimation in such cases typically requires millions of paths, rendering crude Monte Carlo impractical [5, 9].
@@
Classical variance reduction techniques attempt to address these challenges. Antithetic variates reduce variance through negative correlation between paired samples [29, 27], but the theoretical gain is bounded by a factor of two [37]. Control variates can be more effective but require auxiliary variables that are both analytically tractable and highly correlated with the target payoff [46]. For discontinuous payoffs such as binary options, such variables are difficult to construct, limiting applicability.
@@
More advanced methods have been developed. Importance sampling modifies the probability measure to increase the frequency of rare outcomes and applies likelihood ratio weighting to remove bias [27, 26]. Its effectiveness depends on the design of suitable distributions, which is problem-specific and difficult to generalize [51]. Another prominent approach is multilevel Monte Carlo (MLMC), which reduces complexity by combining simulations on coarse and fine discretizations [25, 24]. While efficient for path-dependent derivatives, MLMC is not tailored to extreme event pricing, focusing instead on reducing overall cost.
@@
Recent research combines these techniques to overcome individual limitations. Hybrid methods integrate MLMC with importance sampling to improve efficiency while concentrating sampling in critical regions [1, 36]. Machine learning further enhances importance sampling, with neural networks learning tilting parameters [42] and tensor train decompositions enabling high-dimensional distribution approximation [16].
@@
This work addresses the computational challenges of binary option pricing by applying the adaptive multilevel splitting (AMS) method [Cérou27022007]. AMS extends classical splitting techniques for rare event simulation [21] and builds on the foundations of sequential Monte Carlo [17]. Originally developed in reliability analysis and statistical physics [13, 3, 33, 40], AMS decomposes a rare event into a sequence of more frequent conditional events, transforming a single intractable estimation into multiple tractable subproblems. Although AMS has achieved strong results in other scientific domains, no prior applications are documented in financial derivatives pricing. Recent advances provide theoretical guarantees of unbiasedness and convergence [8, 14, 15], creating a rigorous basis for its use in finance.
@@
Within quantitative finance, interacting particle systems and Sequential Monte Carlo (SMC) methods have already been applied to derivative pricing [34, 45, 47]. Adaptive Multilevel Splitting (AMS) can be understood as a direct evolution of this idea [13]: instead of resampling based on fixed survival sets or generic particle weights, AMS works with a scalar importance function on the whole trajectory and introduces an adaptive sequence of levels determined by the empirical distribution of the particles. At each iteration, a fixed fraction of the least-performing particles is killed and replaced by clones of better-performing ones, so that the intermediate levels separating typical from rare trajectories are constructed on the fly.
@@
The contributions of this study are fourfold. First, an AMS adaptation is introduced for binary option pricing under Black–Scholes and Heston dynamics [6, 30], addressing the specific challenges of risk-neutral valuation and financial time series. Second, the sensitivity of the estimator to parameter choices, including the number of trajectories and resampling rates, is analyzed. Third, numerical experiments compare AMS to standard Monte Carlo, showing substantial gains for deep out-of-the-money options. Fourth, an open source Rcpp implementation is released, supporting Euler, Milstein, and Andersen discretizations [4, 31, 2], two importance functions, and six binary option variants, offering a flexible toolkit for rare event simulation in derivatives pricing. A fully documented and publicly available implementation is provided in the R package amsSim [28], which has been released on CRAN and includes all the algorithms and numerical experiments presented in this paper.
@@ 2.1 Stochastic differential equation models
Numerical experiments are conducted under two standard models for asset price dynamics: the Black–Scholes model [6] and the Heston model [30]. These frameworks allow assessment of the robustness of the AMS approach across different model complexities.
For the Black–Scholes case the exact solution, obtained via logarithmic transformation, removes discretization error [7, 27]:
@@
For the Heston model the variance process requires a scheme that preserves positivity and avoids bias. The quadratic–exponential (QE) method of Andersen [2] is employed, the standard approach for accurate Heston simulation. It matches the first two conditional moments of $V_{t+\Delta t}\,|\,V_{t}$ and selects the update regime according to
@@
with $\epsilon\sim\mathcal{N}(0,1)$. The coefficients $\{K_{0},\dots,K_{4}\}$ and the parameters $a$, $b$, and $\psi$ are given explicitly in [2].
@@ 2.3 Adaptive multilevel splitting (AMS)
Adaptive multilevel splitting (AMS) [Cérou27022007, 13] is a variance reduction method for estimating the probability of rare events. Instead of relying on brute–force Monte Carlo, which wastes almost all trajectories on paths that never approach the rare set, AMS iteratively concentrates simulation effort on trajectories that make progress towards the event of interest.
@@
which is sufficient to preserve unbiasedness [8]. Equality between the two sets is not required, although a score function that is more closely aligned with $D$ leads to a lower variance. For a complete trajectory $X=(X_{t})_{t\in[0,\tau_{f}]}$ we then define its score as
@@ 2.3.1 Theoretical properties of AMS
Let $X=(X_{t})_{t\geq 0}$ be a Markov process with importance function $\xi$ and rare set $D$. For fixed population size $N$ and killing parameter $K\in\{1,\dots,N-1\}$, AMS is well posed: at each iteration the cutting level $Z$ is an order statistic $Z=S^{(K)}$, and under standard regularity assumptions on $X$ (Feller property), on $\xi$ (continuity), and on the entrance into $D$ (strict entrance condition), the algorithm terminates almost surely after a finite number of iterations [13].
@@
without requiring the stronger equivalence $\xi(x)\geq L_{\max}\iff x\in D$ [8, 39, 13]. Unbiasedness extends to unnormalised measures
@@
where the asymptotic variance $\sigma^{2}(\varphi)$ can be characterised via the associated Fleming–Viot particle system [13, 12]. In particular, for $K=1$ and a rare event with probability $p$,
@@
While the unbiasedness of $\hat{p}_{\mathrm{AMS}}$ does not depend on the particular choice of the importance function $\xi$, the variance of the estimator is strongly affected by it. In particular, if $\xi$ is poorly adapted to the geometry of the rare event (for example, in analogy with a badly chosen importance sampling density), the resulting estimator may exhibit large variance and therefore a much slower convergence to the true value. In practice, variance is controlled by experimentally comparing alternative definitions of $\xi$ and by suitably adjusting the population size $N$ and the killing parameter $K$ [13, 8].
@@
AMS adapts intermediate levels and branching rates on the fly, removing the need for a priori specification as in classical Multilevel Splitting [35] or Sequential Monte Carlo [20, 11]. The algorithm maintains a fixed population size $N$, ensuring robustness, parallel efficiency, and predictable memory use. It provides unbiased estimators for both rare event probabilities and unnormalised measures $\gamma(\varphi)$, enabling straightforward parallelization across independent runs [13, 8, 39].
@@ 3.1 Antithetic variates: overview and limitations
Antithetic variates [23] reduce variance by pairing negatively correlated samples. In option pricing this corresponds to simulating each path together with its reflection obtained by negating Brownian increments. For monotone payoffs the estimator variance decreases, with a theoretical maximum reduction by a factor of two.
@@ 3.2 Control variates: overview and limitations
In practice, however, there is no unique, model–independent choice of $W$ that works well across products: an effective control variate must be engineered to closely mimic the structure of the specific payoff under the given dynamics. For example, for arithmetic Asian options one can use the optimized lower-bound proxy of [19]. This makes the technique very powerful when specialized, but at the cost of poor generalizability.
@@ 3.3 Multilevel Monte Carlo: overview and limitations
Multilevel Monte Carlo (MLMC) [25, 24] exploits a hierarchy of approximations $X_{0},\dots,X_{L}$ of the same quantity. The telescoping identity
@@
reduces variance by coupling successive levels with shared randomness. The resulting estimator achieves mean square error $\mathcal{O}(\varepsilon^{2})$ at cost $\mathcal{O}(\varepsilon^{-2})$, compared to $\mathcal{O}(\varepsilon^{-3})$ for standard Monte Carlo [25].
@@ 3.4 Importance sampling: overview and limitations
Importance sampling (IS) [52] estimates $\psi=\mathbb{E}[h(X)]$ by sampling from an alternative density $g$ and reweighting:
@@
In rare event regimes IS becomes unstable. When exercise probabilities are of order $10^{-6}$, the equation $\psi^{\prime}(\theta)=a$ may lack a solution or yield extreme $\theta^{*}$, and evaluation of $e^{\theta Y}$ produces flat likelihood landscapes with sporadic spikes. In such cases Newton–Raphson and related solvers fail to converge, and stochastic optimisers are equally unreliable [10].
@@
Variance explosion: an inappropriate choice of $g(y)$ can inflate the estimator’s variance instead of reducing it [52].
Payoff specific design: effective importance sampling must be tailored to the payoff. Binary calls, binary puts, and Asian options require distinct tilting schemes, and multi-asset payoffs add combinatorial complexity [26].
@@ 4.1 Importance function design
of [8, Theorem 3.2] holds. Then the AMS estimator of the risk–neutral probability $p=\mathbb{Q}(D)$ is
@@ 4.1.1 Heuristics and diagnostics for importance function design
From a theoretical perspective, a first requirement is that $\xi$ preserves the support of the event, in the sense that $x\in D\;\;\Longrightarrow\;\;\xi(x)\geq L_{\max},$ so that the sufficient condition of [8, Theorem 3.2] is satisfied. If this condition were violated, AMS could in principle miss part of $D$ and introduce bias. Even when unbiasedness holds, a poorly informative choice of $\xi$ (for instance, a score that is nearly uncorrelated with $\mathbf{1}_{D}$ or that does not increase as trajectories move towards $D$) inflates the variance and can lead to worse efficiency than crude Monte Carlo.
@@ 4.2 Impact of the selection parameter $K$ on algorithm performance
The selection parameter can be expressed either as the absolute number of discarded trajectories $K$, or equivalently as the discard fraction $k=K/N$, where $N$ is the population size. At each iteration exactly $K=kN$ particles (i.e., a fraction $k$ of the population) are killed and replaced by clones. Theoretical results show that, for an optimal importance function, the asymptotic variance is minimised in the limit $K=1$ [13, 12], but this regime is computationally impractical in realistic settings.
@@ 4.3 Impact of the number of trajectories N on algorithm performance
The particle count $N$ directly affects AMS performance. Larger $N$ reduces estimator variance but increases runtime due to higher simulation and sorting costs. Theoretical analysis shows complexity of order $N(-\log(p))\log(N)$, accounting for the sorting step and the generation of one new trajectory per iteration [13, 12].